A Bibliography of Publications of Lawrence Charles Cowsar

Lawrence Charles Cowsar AT&T Bell Laboratories

600 Mountain Ave., 2C-302A Murray Hill, NJ 07974-0636

USA

Tel: +1 908 582 5468 FAX: +1 908 582 5857

E-mail: cowsar@research.bell-labs.com (Internet)

13 October 2017 Version 1.14

Abstract

This bibliography records publications of Lawrence Charles Cowsar.

Title word cross-reference

1993 [MMM93]. 1997 [BCM01].

Absorbing [CDW96, CDW93]. Adaptive [LSS^{+}98]. Administration [MMM93]. Aeronautics [MMM93]. al [MMM93]. Algorithms [CWW92]. Alto [BCM01]. annual [BCM01]. Applications [CDG^{+}92, Mor93]. approach [KCG^{+}01]. Approximations [CDW96]. April [MMM93].

Balancing [CMW92, CMW93, CMW95]. Boundary [CDW96, CDW93].

CA [BCM01]. Cataluña [Mor93]. Colorado [MMM93]. Computing [CDG^{+}92, DKM^{+}92]. Conditions [CDW96, CDW93]. Conference [DKM^{+}92, MMM93, Mor93]. control [HSF^{+}05]. Copper [MMM93]. cosponsored [MMM93].

Decomposition [CW91, CMW92, CDG^{+}92, CWW92, CMW93, Cow93a, Cow94, GKM^{+}91, CMW95]. Differential [CW91, GKM^{+}91]. dimensional [CRW^{+}93]. Domain [CW91, CMW92, CWW92, CMW93, Cow93a, Cow94, GKM^{+}91, CMW95]. Dual [Cow93b]. Dual-Variable [Cow93b]. dynamics [HSF^{+}05].

Element [ACWY00, Cow93a, CDW96, CDW90, CDW93]. Elements [CW91, CMW93, Cow94, Cow93b, CMW95, Mor93]. Elliptic [CW91, CWW92]. Equations [CW91, CWW92, CDW96, GKM^{+}91, CDW90, CDW93]. Estimates [CDW96, CDW90, CDW93]. evolution [HSF^{+}05].

Fifth [DKM^{+}92]. Finite [ACWY00, CW91, CMW93, Cow93a, Cow94, CDW96, Mor93, CDW90, Cow93b, CDW93, CMW95, Mor93]. flow [CRW^{+}93, KCG^{+}01]. fluids [Mor93]. Fourth [GKM^{+}91].

global [Cow07]. Grids [ACWY00].

held [MMM93, Mor93]. Houston [DKM^{+}92]. Hybrid [Cow94]. Hyperbolic [CDW96, CDW93].

International [GKM^{+}91, Mor93].

Lagrange [Cow93a]. Lagrange-Type [Cow93a].

MAPS [LSS^{+}98]. March [DKM^{+}92]. Mathematical [BCM01]. meeting [BCM01]. Method [CW91]. Methods [ACWY00, Cow93a, Cow93b, GKM^{+}91, MMM93, CDW90, CDW93]. minisymposium [BCM01]. Mixed [ACWY00, CW91, CMW92, CMW93, Cow94, CDW96, CDW90, Cow93b, CDW93, CMW95]. modeling [BCM01, CRW^{+}93]. Mountain [MMM93]. Multiblock [ACWY00]. Multigrid [CWW92, Cow94, MMM93]. multiphase [KCG^{+}01]. Multiresolution [LSS^{+}98].

National [MMM93]. Nonconforming [Cow93a]. Nonmatching [ACWY00].

Oil [CMW92]. optical [BCM01, HSF^{+}05]. optics [BCM01]. optimization [KCG^{+}01]. Order [CDW96, CDW93].

Palo [BCM01]. papers [Mor93]. Parallel [CW91, CWW92, DKM^{+}92]. Parameterization [LSS^{+}98]. Partial [CW91, GKM^{+}91]. Politécnica [Mor93]. power [HSF^{+}05]. Preconditioners [Cow94]. presence [Cow07]. presented [Mor93]. Principles [CDG^{+}92]. priori [CDW90, CDW93, CDW96]. Problems [CMW92]. Proceedings [BCM01, MMM93, DKM^{+}92]. Processing [DKM^{+}92].

research [Cow07]. Reservoir [CMW92].

Schwarz [Cow93b]. science [BCM01]. Scientific [CDG^{+}92, DKM^{+}92]. Second [CDW96, CDW93]. Second-Order [CDW96]. sedimentation [CRW^{+}93]. September [Mor93]. SIAM [BCM01, DKM^{+}92]. Simulation [CMW92, HSF^{+}05]. Sixth [MMM93]. Some [Cow94]. Space [MMM93]. Spaces [Cow93a]. Stanford [BCM01]. Surfaces [LSS^{+}98]. Symposium [GKM^{+}91]. systems [HSF^{+}05].

Their [CDG^{+}92]. transport [HSF^{+}05].

trends [Mor93]. Two [CRW^{+}93]. Two-dimensional [CRW^{+}93]. TX [DKM^{+}92]. Type [Cow93a].

Univ [BCM01]. Universidad [Mor93].

USA [BCM01].

Variable [Cow93b]. VIII [Mor93].

wave [CDW90]. workshop [MMM93].

References

Arbogast:2000:MFE

[ACWY00] Todd Arbogast, Lawrence C. Cowsar, Mary F. Wheeler, and Ivan Yotov. Mixed finite element methods on nonmatching multiblock grids. SIAM Journal on Numerical Analysis, 37(4): 1295–1315, August 2000. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/30844.

Bao:2001:MMO

[BCM01] Gang Bao, Lawrence Cowsar, and Wen Masters, editors. Mathematical modeling in optical science. Proceedings of a minisymposium on optics at SIAM’s annual meeting, Stanford Univ., Palo Alto, CA, USA, 1997, Frontiers in Applied Mathematics. SIAM, Philadelphia, PA, USA, 2001.

Cowsar:1992:DPT

[CDG^{+}92] L. C. Cowsar, E. J. Dean, R. Glowinski, P. Le Tallec, C. H. Li, J. Périaux, and M. F. Wheeler. Decomposition principles and their applications in scientific computing. In Dongarra et al. [DKM^{+}92], pages 213– 237. ISBN 0-89871-303-X. LCCN QA76.58 .P76 1992.

Cowsar:1990:PEM

[CDW90] Lawrence C. Cowsar, Todd F.

Dupont, and Mary F. Wheeler. A priori estimates for mixed finite element methods for the wave equations. Computer Methods in Applied Mechanics and Engineering, 82(1-3):205–222, 1990. CODEN CMMECC. ISSN 0045-7825, 0374-2830. Reliability in computational mechanics (Austin, TX, 1989).

Cowsar:1993:PEM

[CDW93] Lawrence C. Cowsar, Todd F. Dupont, and Mary F. Wheeler. A priori estimates for mixed finite element methods for second order hyperbolic equations with absorbing boundary conditions. Dept. of Comp. and Appl. Math. TR93-45, Rice University, Houston, TX 77251, USA, October

                                        1993. URL gopher://bellona.
                                        cs.rice.edu/11/softlib/CRPC-
                                        TRs/reports/CRPC-TR94464-S.

ps. Published in [CDW96].

Cowsar:1996:PEM

[CDW96] Lawrence C. Cowsar, Todd F. Dupont, and Mary F. Wheeler. A Priori estimates for mixed finite element approximations of second-order hyperbolic equations with absorbing boundary conditions. SIAM Journal on Numerical Analysis, 33(2):492–504, April 1996. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cowsar:1992:BDD

[CMW92] Lawrence Cowsar, Jan Mandel, and Mary F. Wheeler. Balancing domain decomposition for mixed problems in oil reservoir simulation. Presented at the 6th International Symposium on Domain Decomposition Methods, Como, Italy, 1992, 1992.

Cowsar:1993:BDD

[CMW93] Lawrence C. Cowsar, Jan Mandel, and Mary F. Wheeler. Balancing domain decomposition for mixed finite elements.

Dept. Comp. and Appl. Math. TR93-08, Rice University, Houston, TX 77251, USA, March

1993. URL gopher://bellona.
cs.rice.edu/11/softlib/CRPC-
TRs/reports/CRPC-TR93454.ps.

Published in [CMW95].

Cowsar:1995:BDD

[CMW95] Lawrence C. Cowsar, Jan Mandel, and Mary F. Wheeler. Balancing domain decomposition for mixed finite elements. Mathematics of Computation, 64(211):989– 1015, July 1995. CODEN MCM-PAF. ISSN 0025-5718 (paper), 1088-6842 (electronic).

Cowsar:1993:DDMa

[Cow93a] Lawrence C. Cowsar. Domain decomposition methods for nonconforming finite element spaces of Lagrange-type. In Melson et al. [MMM93], pages 93–109. ISBN ???? ISSN 0191-7811. LCCN ???? Two volumes.

Cowsar:1993:DVS

[Cow93b] Lawrence C. Cowsar. Dualvariable Schwarz methods for mixed finite elements. Dept. Comp. and Appl. Math. TR93-09, Rice University, Houston, TX 77251, USA, March 1993. Submitted.

Cowsar:1994:SDD

[Cow94] Lawrence C. Cowsar. Some Domain Decomposition and Multigrid Preconditioners for Hybrid Mixed Finite Elements. PhD thesis, Dept. Comp. Appl. Math., Rice University, Houston, TX 77251-1892, April

                                        1994. URL gopher://bellona.
                                        cs.rice.edu/11/softlib/CRPC-
                                        TRs/reports/CRPC-TR94465-S.
                                        ps.

Cowsar:2007:GRP

[Cow07] Lawrence Cowsar. A global research presence. Bell Labs Technical Journal, 12(2):1–2, Summer 2007. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Chippada:1993:TDM

[CRW^{+}93] S. Chippada, B. Ramaswamy, M. F. Wheeler, L. C. Cowsar, and D. M. Tetzlaff. Two-dimensional modeling of flow and sedimentation. In Morgan [Mor93]. ISBN 84-87867-29-4. LCCN ????

Cowsar:1991:PDD

[CW91] Lawrence C. Cowsar and Mary F. Wheeler. Parallel domain decomposition method for mixed finite elements for elliptic partial differential equations. In Glowinski et al. [GKM^{+}91], pages 358– 372. ISBN 0-89871-278-5. LCCN QA402.2 .I57 199.

Cowsar:1992:PMD

[CWW92] Lawrence C. Cowsar, Alan Weiser, and Mary F. Wheeler. Parallel multigrid and domain decomposition algorithms for elliptic equations. In David E. Keyes, Tony F. Chan, Gérard Meurant, Jeffrey S. Scroggs, and Robert G. Voigt, editors, Fifth International Symposium on Domain Decomposition Methods for Partial Differential Equations (Norfolk, VA, 1991), pages 376–385. SIAM, Philadelphia, PA, USA, 1992. ISBN 0-89871-288-2. LCCN QA402.2 .I57 1991.

Dongarra:1992:PFS

[DKM^{+}92] Jack Dongarra, Ken Kennedy, Paul Messina, Danny C. Sorensen, and Robert G. Voigt, editors. Proceedings of the Fifth SIAM Conference on Parallel Processing for Scientific Computing (Houston, TX, March 25–27, 1991). SIAM, Philadelphia, PA, USA, 1992. ISBN 0-89871-303-X. LCCN QA76.58 .P76 1992.

Glowinski:1991:FIS

[GKM^{+}91] Roland Glowinski, Yuri A.

Kuznetsov, Gérard A. Meurant, Jacques Périaux, and Olof Widlund, editors. Fourth International Symposium on Domain Decomposition Methods for Partial Differential Equations. SIAM, Philadelphia, PA, USA, 1991. ISBN 0-89871-278-5. LCCN QA402.2 .I57 199.

Ho:2005:SPE

[HSF^{+}05] Tin Kam Ho, Todd Salamon, Roland W. Freund, Christopher A. White, Bruce K. Hillyer, Lawrence C. Cowsar, Carl J. Nuzman, and Daniel C. Kilper. Simulation of power evolution and control dynamics in optical transport systems. Bell Labs Technical Journal, 10(1):119–137, Spring 2005. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kearsley:2001:NOA

[KCG^{+}01] A. J. Kearsley, L. C. Cowsar, R. Glowinski, M. F. Wheeler, and I. Yotov. New optimization approach to multiphase flow. Journal of Optimization Theory and Applications, 111(3):473–488, 2001. CODEN JOTABN. ISSN 0022-3239.

Lee:1998:MMA

[LSS^{+}98] Aaron W. F. Lee, Wim Sweldens, Peter Schröder, Lawrence Cowsar, and David Dobkin. MAPS: Multiresolution adaptive parameterization of surfaces. Computer Graphics, 32(Annual Conference Series):95–104, August 1998. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic). URL http:

//w3imagis.imag.fr/Membres/
Frederic.Durand/Book/sig98.
html;   http://www.acm.org:
80/pubs/citations/proceedings/
graph/280814/p95-lee/;   http:
//www.cs.princeton.edu/~wailee/
maps.html.

Melson:1993:SCM

[MMM93] N. Duane Melson, T. A. Manteuffel, and S. F. McCormick, editors. Sixth Copper Mountain Conference on Multigrid Methods: proceedings of a workshop cosponsored by the National Aeronautics and Space Administration ... [et al.] and held at Copper Mountain, Colorado, April 4–9, 1993, number NSAS CP-3224 in NASA conference publication. National Aeronautics and Space Administration, Office of Management, Scientific and Technical Information Program, Washington, DC, USA, 1993. ISBN ???? ISSN 0191-7811. LCCN ???? Two volumes.

Morgan:1993:FEF

[Mor93] K. Morgan, editor. Finite elements in fluids: new trends and applications; [papers presented at the VIII International Conference on Finite Elements in Fluids held at the Universidad Politécnica of Cataluña on 20–23 September 1993]. Centro Internacional de métodos Numéricos en ingeniería, Barcelona, Spain, 1993. ISBN 84-87867-29-4. LCCN ????